A man weighing 80 kg, stands on a weighing scale in a lift which is moving upwards with a uniform acceleration of 5m/s2. What would be the reading on the scale? (g = 10 m/s2)


        (a)    1200 N


        (b)    Zero


        (c)    400 N


        (d)    800 N


Grade:upto college level

2 Answers

Simran Bhatia
10 years ago

        (a)

Reading of the scale

= Apparent wt. of the mass

= m(g + a)

= 80(10 + 5) = 1200 N

Het Patel
4 years ago
In above question it is given that
M=80kg
A=5m/s2
Since we know that weighing scale measures N/G (where N is normal and G is gravitational constant I.e 10 
Now we know that mass and acc. Of lift as well as acc. Due to gravity 
Therefore normal on weighing scale =
N= force exerted by gravity+ force due to acc.
=> N = MG + MA 
=> N = 80(10)+80(5)
=> N = 800+400
=> N = 1200
And that our answer 1200 N